OpenStudy (anonymous):

Find the sum of the first 17 terms of the arithmetic sequence 10, 14, 18, 22, 26...

OpenStudy (anonymous):

arithmetic sequence: sum of the first n term = n/2 (2(1st term) + (n - 1)(difference)) for the sequence: 10, 14, 18, 22, 26, ... the first term is 10 and the difference is 4 sum of the first 17 term = 17/2 (2(10) + (17-1)4) = 17/2 (20 + 64) = 17/2 (84) = 17 (42) = 714